# A phenomenological gravitational waveform model of binary black holes incorporating horizon fluxes

### v2

@@ Acknowledgments
SM thanks the University Grants Commission (UGC), India, for providing financial support as a senior research fellowship. The high performance computing facility Pegasus at IUCAA was used for this work. The authors thank Vaishak Prasad, Apratim Ganguly, Anuj Mishra, Mark Scheel, Nathan Johnson-McDaniel, P. Ajith, and Shrobana Ghosh for useful comments and suggestions. The Python software packages PyCBC [79], Scipy [80], Numpy [81], Matplotlib [82] were used for computational works, analyses, and plots. Mathematica [83] was used for data fitting purposes. SB acknowledges the hospitality of IUCAA with gratitude and support from the NSF under Grant PHY-2309352.
@@ Appendix A Post-Newtonian inspiral phase
Appendix B Ringdown and peak frequency
@@
Appendix C Coefficients for the pseudo-PN parameters
$\sigma_{1}$
$\sigma_{2}$
$\sigma_{3}$
$\sigma_{4}$
$\lambda_{00}$
1650.156
14056.472
--86832.697
146221.791
$\lambda_{10}$
--3531.389
--75378.11
412796.518
--667838.551
$\lambda_{01}$
--2661.071
137530.39
--623322.321
920614.957
$\lambda_{11}$
15385.008
--944130.978
4.1416×$\times 10^{6}$
--5.959×$\times 10^{6}$
$\lambda_{21}$
--219.668
1.7166×$\times 10^{6}$
--7.1273×$\times 10^{6}$
9.8223×$\times 10^{6}$
$\lambda_{02}$
--6845.194
169601.655
--737033.591
1.0493×$\times 10^{6}$
$\lambda_{12}$
38637.011
--1.2802×$\times 10^{6}$
5.4343×$\times 10^{6}$
--7.5726×$\times 10^{6}$
$\lambda_{22}$
--45127.308
2.648×$\times 10^{6}$
--1.09142×$\times 10^{7}$
1.4893×$\times 10^{7}$
$\lambda_{03}$
--2076.335
50611.846
--213200.291
288184.3
$\lambda_{13}$
17433.08
--468141.986
1.9402×$\times 10^{6}$
--2.641×$\times 10^{6}$
$\lambda_{23}$
--31696.712
1.10645×$\times 10^{6}$
--4.5663×$\times 10^{6}$
6.2489×$\times 10^{6}$

### v3

@@ Acknowledgments
SM thanks the University Grants Commission (UGC), India, for providing financial support as a senior research fellowship. KSP gratefully acknowledges the support of the Science and Technology Facilities Council (STFC) of the United Kingdom under the grant ST/V005677/1. The high performance computing facilities Pegasus at IUCAA and Bluebear at the University of Birmingham were used for this work. The authors are grateful for computational resources provided by the LIGO Lab and supported by National Science Foundation Grants PHY-0757058 and PHY-0823459. The authors thank Vaishak Prasad, Apratim Ganguly, Anuj Mishra, Mark Scheel, Nathan Johnson-McDaniel, P. Ajith, and Shrobana Ghosh for useful comments and suggestions. The Python software packages PyCBC [86], Scipy [87], Numpy [88], Matplotlib [89] were used for computational works, analyses, and plots. Mathematica [90] was used for data fitting purposes. SB acknowledges the hospitality of IUCAA with gratitude and support from the NSF under Grant PHY-2309352.
@@ Appendix A Post-Newtonian inspiral phase
Appendix B Nomenclature of various waveform models (existing or temporarily introduced)
Name of the model Description
SEOBNRv2 The EOB waveform of the same name as published in Ref. [16].
SEOBv2 The EOB model after removing the NR calibratios from SEOBNRv2.
SEOBv2_Horizon The time domain model after adding the corrections due to horizon fluxes
of black holes to the phase and the amplitude of SEOBv2.
IMRPhenomD
(abbreviated as PhenomD) The phenomenological IMR waveform of the same name as published in Refs. [24, 25].
IMRPhenomD_Horizon The final frequency-domain IMR model of BBHs with the horizon flux
corrections in phase and amplitude, the main objective of this work.
IMRPhenomD_Recal A recalibrated version of IMRPhenomD we created for comparing with
IMRPhenomD_Horizon, since there are differences in the calibration range and the
TaylorF2 phase that have been used in the former model and the current work.
For this, we use SEOBv2 hybrids instead of SEOBv2_Horizon, and leave out
$\Psi_{\rm TH}$ in Eq. (22).
Table 3:
Appendix C Ringdown and peak frequency
@@
Appendix D Coefficients for the pseudo-PN parameters
$\sigma_{1}$ $\sigma_{2}$ $\sigma_{3}$ $\sigma_{4}$
$\lambda_{00}$ 730.759 25127.527 --130492.02 203261.985
$\lambda_{10}$ 2043.816 --146822.249 714488.221 --1.0914×$\times 10^{6}$
$\lambda_{01}$-5997.126 181096.89 --809867.9 1.1859×$\times 10^{6}$
$\lambda_{11}$ 54465.793 --1.48114×$\times 10^{6}$ 6.566×$\times 10^{6}$-9.5763×$\times 10^{6}$
$\lambda_{21}$-88292.028 2.9436×$\times 10^{6}$-1.275×$\times 10^{7}$ 1.8316×$\times 10^{7}$
$\lambda_{02}$-12729.742 249044.619 --1.0889×$\times 10^{6}$ 1.5653×$\times 10^{6}$
$\lambda_{12}$ 118443.267 --2.3723×$\times 10^{6}$ 1.0356×$\times 10^{7}$-1.4902×$\times 10^{7}$
$\lambda_{22}$-275615.32 5.7947×$\times 10^{6}$-2.51×$\times 10^{7}$ 3.6013×$\times 10^{7}$
$\lambda_{03}$-4715.744 86494.178 --373232.315 524127.359
$\lambda_{13}$ 52582.588 --947035.327 4.0911×$\times 10^{6}$-5.8335×$\times 10^{6}$
$\lambda_{23}$-139301.918 2.5658×$\times 10^{6}$-1.1108×$\times 10^{7}$ 1.5942×$\times 10^{7}$